OpenStudy (anonymous):

Another aquarium in the building has dimensions that are each triple the size of the aquarium in Riley’s father’s office. Riley thought that the volume would also triple. Is Riley correct? How many times greater is the volume of the larger aquarium than the smaller? explain your reasoning. help please

OpenStudy (anonymous):

@KamiBug

OpenStudy (kamibug):

The original dimensions were 16, 8.5, and 10.5, correct? :)

OpenStudy (anonymous):

yeah forogt to put it

OpenStudy (anonymous):

wait let me put it

OpenStudy (anonymous):

Riley noticed an aquarium in his father’s office. The aquarium had the dimensions 16 in. by 8.5 in. by 10.5 in. The formula for volume is: V = l × w × h.

OpenStudy (cwrw238):

the ratio of the volumes equals the ratio of the cubes of the lengths

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait what?

OpenStudy (cwrw238):

so if one dimension is 10 then the corresponding dimension in the larger aquarium will be 30 and ratio of these is 1:3 so ratio of the volumes will be 1 : 3^3 = 1:27

OpenStudy (anonymous):

so my answer is 1:27 sir??

OpenStudy (kamibug):

Multiply each original dimension by 3 and that'll give you the new dimensions. Use those new dimensions to solve for the new volume. Then divide the original volume by the new volume to see how much bigger it is! :D 16*3=48 8.5*3=25.5 10.5*3=31.5 48*25.5*31.5=38556 Now divide the new volume by the original volume... 38556/1428 = 27 So the bigger aquarium is 27 times bigger than the smaller one! :)

OpenStudy (anonymous):

so the answer is 27?

OpenStudy (anonymous):

this makes way way more sense :)

OpenStudy (anonymous):

so i times 27 and 3?

OpenStudy (kamibug):

You don't need to multiply 27 and 3. :P The work is already done, Lol. Riley was incorrect b/c he thought it would only be 3 times bigger; it's actually 27 times bigger. :) Just explain that when you put your answer. :3